OpenStudy (loser66):

I have to use Gauss's lemma to show that if \(P(x) = a_nx^n+....+a_1x+a_0\) in \(\mathbb Z[x]\) has a rational root r/s, where gcd (r,s) =1, then \(s|a_n\) and \(r|a_0\) My problem is: I can prove it without using Gauss's lemma. How to link my proof to Gauss's lemma? Please, help My proof is in comment box

OpenStudy (loser66):

Suppose r/s is a rational root of P(x) , then P(r/s)=0 \(a_n(r/s)^n+......+a_1(r/s)+a_0=0\\a_0= -(a_n(r/s)^n+.....+a_1r/s\\a_0=r/s(-a_n(r/s)^{n-1}-....-a_1)\) the second term is a polynomial but \(a_0\) is a number, hence only 1 possibility is r/s | a_0

OpenStudy (loser66):

\(r/s| a_0\) --> \(r = a_0 s\) that is \(r|a_0\)

OpenStudy (rational):

Basically you want to prove rational root theorem using Gauss lemma

OpenStudy (rational):

whats Gauss's lemma anyways ?

OpenStudy (loser66):

Gauss's lemma: Let R be a UFD with field of fractions F a) The product of 2 primitive elements of R[x] is primitive b) Suppose f(x) in R[x] .Then f(x) has a factorization f(x) = g(x) h(x) in F[x] with degree of g, h greater or equal 1 if and only if f(x) has such a factorization in R[x]

OpenStudy (loser66):

We can't use part a of lemma here, since we don't know whether P(x) is primitive or not.

OpenStudy (rational):

I found this http://en.wikipedia.org/wiki/Rational_root_theorem they're dividing gcd of the coefficients of polynomial to make it primitive

OpenStudy (loser66):

Thank you so much. It is waaaay better than my proof. :)